# Multistability of graphene nanobubbles

### v1

@@ V Conclusion
Using the example of inert gas atoms (Ar, Kr, Xe), it has been shown that graphene nanobubbles located on flat substrates are multistable systems. A nanobubble can have many stable stationary states, each characterized by the number of layers $l$ in the internal atom cluster. Inside the nanobubble, each atom lies within one of the layers. The layers have a circular, flat shape, are concentrically stacked on top of each other, and form an $l$-stepped pyramid with a flat top. Atoms within each layer lie in a single plane, with only the edge atoms displaced downwards, smoothing the stepped shape of the pyramid. The covering of the pyramid from above by a graphene sheet occurs due to its local stretching. The sheet’s valence bonds are stretched only above the group of encapsulated atoms; outside this coverage zone, the sheet remains undeformed and tightly adheres to the substrate. The strongest stretching (up to one percent) occurs above the edges of the layers.
The maximum possible number of layers $l_{m}$ increases monotonically with the number of internal atoms $N_{g}$. For argon atoms, only single-layer stationary states ($l_{m}=1$) are possible for $N_{g}\leq 150$; for $150<N_{g}<900$—single-layer and double-layer states ($l_{m}=2$); for $900<N_{g}<1800$, $l_{m}=3$; for $1800<N_{g}<2800$, $l_{m}=4$; and for $2800<N_{g}<4000$, $l_{m}=5$. The graphene sheet, due to van der Waals interaction with the substrate, presses the cluster of encapsulated atoms to substrate, creating pressures $P\sim 1$ GPa. The highest pressure is experienced by single-layer packings of internal atoms. The magnitude of the pressure decreases as the number of layers increases.
Thermal fluctuation modeling shows that among all $l$-layer states of the nanobubble, there is always one ”ground” state, which smoothly transitions into a layerless liquid state as the temperature increases. The other states transition into this state at a specific temperature $T_{l}$. For example, for $N_{g}=2909$ encapsulated atoms, the ground state is the four-layer packing ($l=4$). For Xe atoms, the $l$-layer states persist up to temperatures $T_{l}=960$, 390, 330, 210 K; for Ar atoms, $T_{l}=600$, 270, 180, 30 K; and for Kr atoms—750, 360, 240, 150 K, where $l=1$, 2, 3, 5.
The existence of several stable states with different numbers of layers in the internal atom cluster at low temperatures indicates that universal scaling does not hold for these nanobubbles. The height-to-radius ratio $H/R$ of the nanobubble is not constant and can vary from 0 to 0.24. The universal relation $H/R\approx 0.2$ holds only for the ground states.
@@ Acknowledgements
The research was funded by the Russian Science Foundation (RSF) (project No. 25-73-20038).

### v3

@@ V Conclusion
Using the example of inert gas atoms (He, Ne, Ar, Kr, Xe), it has been shown that graphene nanobubbles on flat substrates exhibit multistability. These bubbles can adopt multiple stable stationary states, each defined by a distinct number of layers, $l$, within an internal cluster of atoms. Within the nanobubble, each atom is confined to one of these layers. The layers themselves are circular, flat, and concentrically stacked, forming an $l$-stepped pyramid with a flat top. While atoms within each layer occupy a single plane, the edge atoms are displaced downward, effectively smoothing the stepped profile of the pyramid. This pyramidal cluster is encapsulated by a locally stretched graphene sheet. The sheet’s valence bonds elongate only directly above the group of confined atoms; outside this coverage zone, the membrane remains undeformed and tightly adhered to the substrate. The most significant elongation, reaching up to one percent, occurs directly above the layer edges.
The maximum number of possible layers $l_{m}$ increases monotonically with the total number of encapsulated atoms, $N_{g}$. For argon atoms, only single-layer stationary states ($l_{m}=1$) are possible for $N_{g}<145$; for $145<N_{g}<550$, both single-layer and double-layer states exist ($l_{m}=2$); for $550<N_{g}<1600$, triple-layer states are added ($l_{m}=3$); for $1600<N_{g}<2500$, four-layer states appear ($l_{m}=4$); and five-layer states ($l_{m}=5$) become accessible for $2500<N_{g}<4000$. The graphene membrane, through van der Waals interaction with the substrate, exerts pressure on the encapsulated atomic cluster, confining it against the substrate with pressures on the order of $P\sim 1$ GPa. This pressure is most pronounced for single-layer atomic packings and systematically decreases as the number of layers increases.
Thermal fluctuation modeling reveals that among all $l$-layer states of a graphene nanobubble, one ”ground” state always exist. This state smoothly transitions into a layerless liquid configuration as the temperature increases, whereas the remaining $l$-layer states undergo a transition into this ground state at a distinct temperature $T_{l}$. For a cluster of $N_{g}=4000$ encapsulated atoms, the ground state corresponds to the four-layer packing ($l=4$). The thermal stability of $l$-layer states varies significantly with the atomic species. For helium atoms, the $l$-layer states persist up to temperatures $T_{l}=290$, $72$, $47$, and $23$ K for $l=1$, 22, 33, and $5$, respectively. For neon, the corresponding stability limits are 390, 120, 90, and 60 K; for argon—630, 270, 180, and 120 K; for krypton—720, 300, 240, and 180 K; and for xenon—780, 390, 330, and 240 K.
The coexistence of multiple stable states with distinct layer numbers at low temperatures demonstrates that universal scaling does not apply to these graphene nanobubbles. Specifically, the height-to-radius ratio $H/R$ is not constant and can vary across the range from 0 to 0.28. The commonly reported universal relation $H/R\approx 0.2$ holds only for the ground states of the system.
@@ Acknowledgements
Computational facilities were provided by the Joint Supercomputer center (JSCC) of the National Research Center ”Kurchatov Institute”. The research was funded by the Russian Science Foundation (RSF) (project No. 25-73-20038).
Data Availability
All data are openly available data